# 4d Maxwell on the Edge: Global Aspects of Boundary Conditions and Duality

### v1

@@ 2.1 Free boundary conditions
$S_{3d}^{N}=\frac{i}{2\pi}\int_{\partial X_{4}}\Phi^{1}\wedge d\Phi^{2}-\Phi^{2}\wedge dA\,.$ (2.5)
Here $\Phi,\Phi^{1}$ and $\Phi^{2}$ are $U(1)$ gauge fields defined on the boundary $\partial X_{4}$. Indeed, by integrating them out, one gets
@@
$\begin{split}&\text{Dirichlet: }\frac{i}{e^{2}}*F-\frac{\theta}{2\pi}F=d\Phi\,,\\ &\text{Neumann: }\frac{i}{e^{2}}*F-\frac{\theta}{2\pi}F=-d\Phi^{2}=0\,.\end{split}$ (2.8)
In fact, an equivalent way to encode Neumann boundary conditions is to have no edge modes at all, which indeed follows by integrating out $\Phi^{2}$ and then $\Phi^{1}$. One can alternatively consider the shift $\Phi^{1}\to\Phi^{1}+A$, which decouples the edge modes from the bulk. It straightforwardly yields the same boundary equations of motion as above.
@@
where $n,m\in\mathbb{Z}$ are respectively the charges of these operators under the electric and magnetic symmetries. [^5] The minus sign in the definition of the ’t Hooft line $H_{m}[\gamma]$ in terms of the dual field $\tilde{A}$ is crucial and it can be derived by explicitly computing the linking phase between this operator and the symmetry operator $U_{\beta}^{m}=\exp{\left(i\beta\int*J_{m}\right)}$. The Dirichlet (resp. Neumann) boundary conditions defined in (2.3) preserves the $U(1)$ global symmetry generated by $J_{e}$ (resp. $J_{m}$), yielding a $U(1)$ 0-form symmetry on the boundary.
@@ 2.1.1 Global properties of the boundary conditions
The crucial subtlety that we want to focus on is that the conditions (2.3) do not fully determine the boundary conditions. While they fix the field strength, they leave the holonomy of the gauge field undetermined—an additional gauge-invariant piece of data associated with the principal bundle. However, this piece of data is completely constrained by the actions (2.4) and (2.5); by summing over the flux sectors $\int d\Phi^{i}\in 2\pi\mathbb{Z}$, we get the conditions
@@
$\text{Neumann: }\quad H_{q}[\gamma]\equiv\exp{\left(-iq\int\Phi^{2}\right)}=1\quad\forall q\,.$ (2.14)
@@
$\text{Neumann: }\quad S_{3d}^{N,\tilde{r}}=\frac{i}{2\pi}\int_{\partial X_{4}}\tilde{r}\Phi^{1}\wedge d\Phi^{2}-\Phi^{2}\wedge dA\,.$ (2.16)
@@
$\text{Neumann: }\int\Phi^{2}\equiv\int\tilde{A}\in\frac{2\pi}{\tilde{r}}\mathbb{Z}\,$, (2.18)
@@ 2.1.2 Generic free boundary conditions
where $p\in\mathbb{Z}$, $\Phi=(\Phi^{1},\cdots,\Phi^{n})^{T}$ is a vector of 1-form $U(1)$ gauge fields living on the 3d boundary, $k_{j\ell}$ ($j,\ell=1,\dots,n$) is a symmetric integral matrix (ensuring proper quantization under large gauge transformations), and $v_{\ell}$ ($\ell=1,\dots,n$) is a vector of integers.
@@
$\int\frac{ip}{4\pi}A\wedge dA=\int\frac{ip}{2\pi}A\wedge d\Phi^{n+1}-\frac{ip}{4\pi}\Phi^{n+1}\wedge d\Phi^{n+1}$ (2.21)
@@
As before, the equations of motion for $\Phi^{\ell}$ impose local constraints on the bulk field strength. Assuming that the matrix $k$ is non-degenerate, [^7] The case of a degenerate $k$, which is related to Dirichlet-like boundary conditions, will be treated at the end of this subsection. we obtain:
@@
Summing over the holonomies of $\Phi^{\ell}$ gives the $n+1$ following constraints:
@@ 3.1 Duality interfaces
$k_{f}=\begin{pmatrix}k&v&0_{n+1}\\ v^{T}&\beta\delta&-\alpha\delta\\ 0_{n+1}^{T}&-\alpha\delta&\alpha\gamma\end{pmatrix}\,,\quad v_{f}=\begin{pmatrix}0_{n+1}\\ \beta\\ -\alpha\end{pmatrix}\,,\quad\Phi^{i}_{f}=\begin{pmatrix}\Phi_{3d}^{i}\\ A\\ \Phi\end{pmatrix}\$, (3.19)
where $0_{n+1}$ is a column vector of $(n+1)$ zeros and $\Phi_{3d}^{i}$ are the $(n+1)$ edge modes of the initial boundary. The new boundary conditions are characterized by
@@ 3.2 Rescaling interfaces
$k_{f}=\begin{pmatrix}k&v&0_{n+1}\\ v^{T}&0&-\mathcal{N}\\ 0^{T}_{n+1}&-\mathcal{N}&0\end{pmatrix}\quad,\quad v_{f}=\begin{pmatrix}0_{n+1}\\ 0\\ -\mathcal{M}\end{pmatrix}\quad,\quad\Phi^{i}_{f}=\begin{pmatrix}\Phi_{3d}^{i}\\ A\\ \Psi\end{pmatrix}$ (3.40)
@@ 3.2.1 Condensation defect
$\hat{k}=\begin{pmatrix}k_{f}&v_{f}&0\\ v_{f}^{T}&0&-\mathcal{M}\\ 0^{T}&-\mathcal{M}&0\end{pmatrix}\quad,\quad\hat{v}=\begin{pmatrix}0\\ 0\\ -\mathcal{N}\end{pmatrix}\,\quad,\quad\hat{\Phi}^{i}=\begin{pmatrix}\Phi^{i}\\ A^{\prime}\\ \Psi^{\prime}\end{pmatrix}=\begin{pmatrix}\Phi_{3d}^{i}\\ A\\ \Psi\\ A^{\prime}\\ \Psi^{\prime}\end{pmatrix}\$. (3.53)
@@ 4 Interacting Boundary Conditions and Boundary CFTs
as the new boundary condition which couples the 3d CFT to the bulk photon. This boundary condition can also be interpreted as defining a non-local 3d CFT, which we denote by $B^{(p,v,k,u)}(\tau,\overline{\tau})$ [32]. This theory arises from integrating out the bulk photon to obtain an effective three-dimensional description. However, due to bulk interactions, the resulting theory is generically non-local and lacks a stress tensor.
@@ 4.1 Weak coupling limit and 3d CFTs
$S_{3d}=S_{CFT}+i\int\Phi^{1}\wedge*J_{3d}+\frac{iN}{2\pi}\int\Phi^{1}\wedge d\Phi^{2}+\frac{i\xi}{4\pi}\int\Phi^{1}\wedge d\Phi^{1}\,.$ (4.11)
@@
$S^{N}_{\partial}$ $=S_{\mathrm{CFT}}+\int\left(\frac{i}{2\pi}\Phi^{1}\wedge d\Phi^{2}+\frac{i}{2\pi}\Phi^{2}\wedge dA-i\Phi^{1}\wedge*J_{3d}\right)=S_{\mathrm{CFT}}+\int iA\wedge*J_{3d}\,$, (4.13)
where in the last equation we have integrated out $\Phi^{2}$. In the weak coupling limit they generate the local CFTs
@@ 5.1 The 5d SymTFT
$\mathcal{B}(U_{x}V_{y}(\Gamma),U_{x^{\prime}}V_{y^{\prime}}(\Gamma^{\prime}))=e^{2\pi i(xy^{\prime}-yx^{\prime})\mathrm{Link}(\Gamma,\Gamma^{\prime})}\,.$ (5.4)

### v2

@@ 2.1 Free boundary conditions
$S_{3d}^{N}=\frac{i}{2\pi}\int_{\partial X_{4}}\Phi_{1}\wedge d\Phi_{2}-\Phi_{2}\wedge dA\,.$ (2.5)
Here $\Phi,\Phi_{1}$ and $\Phi_{2}$ are $U(1)$ gauge fields defined on the boundary $\partial X_{4}$. Indeed, by integrating them out, one gets
@@
$\begin{split}&\text{Dirichlet: }\frac{i}{e^{2}}*F-\frac{\theta}{2\pi}F=d\Phi\,,\\ &\text{Neumann: }\frac{i}{e^{2}}*F-\frac{\theta}{2\pi}F=-d\Phi_{2}=0\,.\end{split}$ (2.8)
In fact, an equivalent way to encode Neumann boundary conditions is to have no edge modes at all, which indeed follows by integrating out $\Phi_{2}$ and then $\Phi_{1}$. One can alternatively consider the shift $\Phi_{1}\to\Phi_{1}+A$, which decouples the edge modes from the bulk. It straightforwardly yields the same boundary equations of motion as above.
@@
where $n,m\in\mathbb{Z}$ are respectively the charges of these operators under the electric and magnetic symmetries. [^5] The minus sign in the definition of the ’t Hooft line $H_{m}[\gamma]$ in terms of the dual field $\tilde{A}$ follows from using the same conventions as in [54]. The Dirichlet (resp. Neumann) boundary conditions defined in (2.3) preserves the $U(1)$ global symmetry generated by $J_{e}$ (resp. $J_{m}$), yielding a $U(1)$ 0-form symmetry on the boundary.
@@ 2.1.1 Global properties of the boundary conditions
The crucial subtlety that we want to focus on is that the conditions (2.3) do not fully determine the boundary conditions. While they fix the field strength, they leave the holonomy of the gauge field undetermined—an additional gauge-invariant piece of data associated with the principal bundle. However, this piece of data is completely constrained by the actions (2.4) and (2.5); by summing over the flux sectors $\int d\Phi_{i}\in 2\pi\mathbb{Z}$, we get the conditions
@@
$\text{Neumann: }\quad H_{q}[\gamma]\equiv\exp{\left(-iq\int\Phi_{2}\right)}=1\quad\forall q\,.$ (2.14)
@@
$\text{Neumann: }\quad S_{3d}^{N,\tilde{r}}=\frac{i}{2\pi}\int_{\partial X_{4}}\tilde{r}\Phi_{1}\wedge d\Phi_{2}-\Phi_{2}\wedge dA\,.$ (2.16)
@@
$\text{Neumann: }\int\Phi_{2}\equiv\int\tilde{A}\in\frac{2\pi}{\tilde{r}}\mathbb{Z}\,$, (2.18)
@@ 2.1.2 Generic free boundary conditions
where $p\in\mathbb{Z}$, $\Phi=(\Phi_{1},\cdots,\Phi_{n})^{T}$ is a vector of 1-form $U(1)$ gauge fields living on the 3d boundary, $k_{j\ell}$ ($j,\ell=1,\dots,n$) is a symmetric integral matrix (ensuring proper quantization under large gauge transformations), and $v_{\ell}$ ($\ell=1,\dots,n$) is a vector of integers.
@@
$\int\frac{ip}{4\pi}A\wedge dA=\int\frac{ip}{2\pi}A\wedge d\Phi_{n+1}-\frac{ip}{4\pi}\Phi_{n+1}\wedge d\Phi_{n+1}$ (2.21)
@@
As before, the equations of motion for $\Phi_{\ell}$ impose local constraints on the bulk field strength. Assuming that the matrix $k$ is non-degenerate, [^7] The case of a degenerate $k$, which is related to Dirichlet-like boundary conditions, will be treated at the end of this subsection. we obtain:
@@
Summing over the holonomies of $\Phi_{\ell}$ gives the $n+1$ following constraints:
@@ 3.1 Duality interfaces
$k_{f}=\begin{pmatrix}k&v&0_{n+1}\\ v^{T}&\beta\delta&-\alpha\delta\\ 0_{n+1}^{T}&-\alpha\delta&\alpha\gamma\end{pmatrix}\,,\quad v_{f}=\begin{pmatrix}0_{n+1}\\ \beta\\ -\alpha\end{pmatrix}\,,\quad\Phi_{f,i}=\begin{pmatrix}\Phi_{i}^{3d}\\ A\\ \Phi\end{pmatrix}\$, (3.19)
where $0_{n+1}$ is a column vector of $(n+1)$ zeros and $\Phi_{i}^{3d}$ are the $(n+1)$ edge modes of the initial boundary. The new boundary conditions are characterized by
@@ 3.2 Rescaling interfaces
$k_{f}=\begin{pmatrix}k&v&0_{n+1}\\ v^{T}&0&-\mathcal{N}\\ 0^{T}_{n+1}&-\mathcal{N}&0\end{pmatrix}\quad,\quad v_{f}=\begin{pmatrix}0_{n+1}\\ 0\\ -\mathcal{M}\end{pmatrix}\quad,\quad\Phi_{f,i}=\begin{pmatrix}\Phi^{3d}_{i}\\ A\\ \Psi\end{pmatrix}$ (3.40)
@@ 3.2.1 Condensation defect
$\hat{k}=\begin{pmatrix}k_{f}&v_{f}&0\\ v_{f}^{T}&0&-\mathcal{M}\\ 0^{T}&-\mathcal{M}&0\end{pmatrix}\quad,\quad\hat{v}=\begin{pmatrix}0\\ 0\\ -\mathcal{N}\end{pmatrix}\,\quad,\quad\hat{\Phi}_{i}=\begin{pmatrix}\Phi_{i}\\ A^{\prime}\\ \Psi^{\prime}\end{pmatrix}=\begin{pmatrix}\Phi_{i}^{3d}\\ A\\ \Psi\\ A^{\prime}\\ \Psi^{\prime}\end{pmatrix}\$. (3.53)
@@ 4 Interacting Boundary Conditions and Boundary CFTs
as the new boundary condition which couples the 3d CFT to the bulk photon. This boundary condition can also be interpreted as defining a non-local 3d CFT, which we denote by $B^{(p,v,k,u)}(\tau,\overline{\tau})$ [34]. This theory arises from integrating out the bulk photon to obtain an effective three-dimensional description. However, due to interactions with the bulk, the resulting theory is generically non-local and lacks a stress tensor. A caveat is that these interactions may induce relevant deformations that spoil boundary conformal symmetry and produce an RG flow. As emphasized in [34], at the perturbative level one can always tune the boundary couplings to cancel such contributions and restore conformal invariance order by order in perturbation theory. However, in some cases non-perturbative effects may drive the boundary theory into a phase in which conformality is lost. If we wish to restrict to conformal boundary conditions, we must therefore assume that such effects are absent and exclude the 3d CFTs for which this occurs.
@@ 4.1 Weak coupling limit and 3d CFTs
$S_{3d}=S_{CFT}+i\int\Phi_{1}\wedge*J_{3d}+\frac{iN}{2\pi}\int\Phi_{1}\wedge d\Phi_{2}+\frac{i\xi}{4\pi}\int\Phi_{1}\wedge d\Phi_{1}\,.$ (4.11)
@@
$S^{N}_{\partial}$ $=S_{\mathrm{CFT}}+\int\left(\frac{i}{2\pi}\Phi_{1}\wedge d\Phi_{2}+\frac{i}{2\pi}\Phi_{2}\wedge dA-i\Phi_{1}\wedge*J_{3d}\right)=S_{\mathrm{CFT}}+\int iA\wedge*J_{3d}\,$, (4.13)
where in the last equation we have integrated out $\Phi_{2}$. In the weak coupling limit they generate the local CFTs
@@ 5.1 The 5d SymTFT
$\langle U_{x}V_{y}(\Gamma)\ U_{x^{\prime}}V_{y^{\prime}}(\Gamma^{\prime})\rangle=e^{2\pi i(xy^{\prime}-yx^{\prime})\mathrm{Link}(\Gamma,\Gamma^{\prime})}\,.$ (5.4)